### v1

@@ 3.1 Mining SO Data
[…]
@@
Questions 674 425674425674\,425674 425 578 115578115578\,115578 115
Questions with Answers 424 808424808424\,808424 808 329 412329412329\,412329 412
@@
212 775212775212\,775212 775 157 187157187157\,187157 187
Questions without Answers 249 617249617249\,617249 617 248 703248703248\,703248 703
Answers 543 533543533543\,533543 533 425 391425391425\,391425 391
Comments 1 673 90616739061\,673\,9061 673 906 1 348 98213489821\,348\,9821 348 982
Tags 36 8373683736\,83736 837 37 3283732837\,32837 328
@@
Initially, we collected the posts and their related comments (as returned by Stack Overflow’ API). However, in Stack Overflow’s data, questions and answers are interchangeably treated as regular posts, differentiating from each other based on their type. Furthermore, while questions can receive multiple answers, only one can be targeted as the accepted one; posts can be associated with multiple comments. We collected all information available for each element under investigation, including the creation and last modification date, title, description, score, views, etc. The creation and modification dates are essential information since a question might have been added before the ChatGPT release, but its accepted answer might have been dated after ChatGPT’s release. Next, we organize the data by associating the questions with their related answers and comments. This step is required since answers and comments for questions not dated from the interval evaluated here are not eligible for our study. For the questions posted before ChatGPT release, we only consider as valid answers and comments, the ones also dated before the release date (30th November 2022) Eventually, we were able to collect 674 425674425674\,425674 425 questions before ChatGPT’s release and 578 115578115578\,115578 115 after (we already notice a drop of almost 100 000100000100\,000100 000 for the same amount of time, i.e., five months). Table 1 presents an overview of our dataset.
@@
To investigate the impact of questions addressing programming languages, frameworks and library issues, we first properly filtered the questions focusing on these subjects. When users ask about these domains, they commonly add the name of the target topic as one of the question tags; since multiple tags can be used for a single question, we just counted the number of times a single tag was adopted. Although users may consider variations or abbreviations for targeting their subjects, we observe a consistent adoption for the target topics evaluated here. This way, we consider their regular official names, like JavaScript, Pandas, and Angular.
@@
Next, based on the final set of questions and related answers and comments, we further explored the associated users. For that, we group them into three groups: questioners, respondents, and commentators. This way, we could compare the number of users involved in each type of activity in Stack Overflow. To answer RQ4, we collected further user information, including their account creation date, leading us to identify reminiscent and new users. We request each user’s information using its ID on the Stack Overflow API.
@@ 3.2 Prompting LLM Models
This study focuses on two LLMs: ChatGPT 3.5 and LLaMA-2. The ChatGPT model, developed by OpenAI, can be accessed via a paid API or a free graphical user interface (GUI). We used the GPT-3.5-turbo variant for our research, interacting with it through its API. On the other hand, LLaMA-2 is a model provided by Meta and is freely accessible for research and commercial purposes. Unlike ChatGPT, LLaMA requires manual loading by the user, as it currently lacks an API, providing a desktop application only for Mac operational systems. For our experiments, we utilized the 7B-sized LLaMA-2 model, which was loaded via Hugging Face support [^5] https://huggingface.co/meta-llama/Llama-2-7b, on a machine equipped with 40GB RAM.
When prompting the LLMs, we follow a standardized approach that simulates a prior conversation to establish context. For that, we consider the tags associated with the questions as the foundation for instructing the LLM to adopt the persona of an expert with extensive knowledge in the relevant subject matterWhite et al. (2023). Next, we consider the question’s title to properly ask the LLM to explain how to fix the problem. Additionally, we argue that further context will be given by informing the question description. Figure 2 presents the adopted approach with more details.
@@
Figure 2: Prompt approach adopted to prompt the LLMs. Questions contents were given as input, replacing the elements informed highlighted with << and >>, respectively.
3.3 Collecting Metrics
@@
In this section, we explain the metrics and tests we consider when running our study. We address each RQ under investigation here, discussing their related methodology steps.
3.3.1 RQ1 and RQ5: General Statistics.
We conducted various statistical analyses on different components of our dataset, including questions, answers, and comments. To assess the impact of ChatGPT’s release on question frequency, we grouped questions by the day they were posted and then analyzed them. A similar methodology was adopted for evaluating answers and comments. Furthermore, given that our data did not follow a normal distribution, as verified by the Shapiro-Wilk test (Shapiro and Wilk, 1965),666p-value < 0.01 we employed the non-parametric Wilcoxon test to compare the distributions of our different datasets (Mann and Whitney, 1947).
3.3.2 RQ2: Similarity Analysis
Textual Similarity. To assess the similarity between answers from Stack Overflow and those generated by LLMs, we employed the cosine similarity metric (Salton and Buckley, 1988; Lahitani et al., 2016). This metric represents a measure for comparing two non-zero vectors ranging from -1 to 1. While a 1 value indicates that the two vectors are identical, a -1 indicates complete dissimilarity, and a 0 value suggests that the vectors have no similarity. To compute this metric, we relied on a pre-trained Sentence Transformers model (all-MiniLM-L6-v2) [^7] https://github.com/UKPLab/sentence-transformers. Both sets of answers were initially embedded into PyTorch tensors [^8] https://pytorch.org/tutorials/beginner/blitz/tensor_tutorial.html, following which the cosine similarity was calculated.
Semantic Similarity. Knowing that the cosine similarity computes the similarity based on how similar the vectors are in their magnitude and direction, not considering the involved semantics, we might misclassify similar answers due to low similarity based on the answers’ structure and style. This way, we also compute the similarity considering the semantics by repeatedly prompting the LLMs, as they might explore word embeddings and pre-trained models like BERT (Devlin et al., 2018). We adopt a similar approach when using the models to answer the questions. First, we simulate a prior conversation to establish the context, using again the questions’ tags to instruct the LLM to adopt the persona of an expert. Next, we ask the LLM to compare the original answer from Stack Overflow and the one generated by the LLM. The expected output adopts a single value based on a VERY LOW to VERY HIGH scale. While ChatGPT did not classify one answer due to the maximum of supported tokens, LLaMA did not classify 72 answers due to resource constraints.
Finally, we also compute Krippendorff’s alpha (Krippendorff, 2011) as a metric for assessing the level of inter-rater agreement between our manual analysis and the assessments provided by LLMs, and the agreements between the LLMs themselves. For the assessment involving our manual analysis, a subsample of 91 questions was randomly selected and manually labeled by the primary author, ensuring a 95% confidence level with a 10% margin of error. Next, we perform the different comparisons based on the available classifications.
@@ 3.3.2 RQ2: Similarity Analysis
Sentiment Analysis. In light of research suggesting that avoiding negative comments can improve answer acceptance on Stack Overflow (Calefato et al., 2015), we aim to investigate whether LLMs’ responses align with this observation. To do so, we conducted a sentiment analysis on answers generated by the LLMs using a pre-trained Transformers model (sentiment-analysis) [^9] https://github.com/huggingface/transformers. Subsequently, we compared the sentiment outcome achieved by each LLM used in our study. We employed the Wilcoxon test for this comparative analysis to identify any statistically significant differences in the reported sentiment levels.
Similarity Correlation. To explore the factors contributing to the reliability of answers generated by LLMs, we extended our analysis to include the context in which the models were prompted. For that, we compute the correlation of the textual similarity, previously computed, with some context information such as questions’ title, description length, and associated tags. Since our dataset was not drawn from a normally distributed population, as confirmed by a Shapiro-Wilk test, we used the non-parametric Spearman test.
(a) Questions
(b) Answers
(c) Comments
Figure 3: Violin plot of posted contents on Stack Overflow.
3.3.3 RQ3: Manual Analysis and Comparison of Generated Answers
To delve deeper into the challenges faced by LLMs when trying to address the Stack Overflow questions, we perform a manual analysis focusing on the questions linked to answers categorized as demonstrating low and high textual similarity for each LLM evaluated here. For that, we curate and analyze the questions associated with the first and fourth quartile of each LLM violin plot (low and high similarity, see Figure 6). Next, the first author of this study analyzed each question, characterizing them based on a set of labels and their associated values. These labels were formulated through an initial exploratory analysis of the questions aimed at distilling the core informational components inherent to each question. Since the main focus here is to report the domains that challenge LLMs, we analyze the questions based on the categories reported by Barua et al. (2014). Some previous related categories were combined together; for example, UI Development and Website Design/CSS was moved together to the category Web Development. A meticulously designed spreadsheet was then established, incorporating these labels and associated values, to provide a systematic guide for the manual analysis.
3.3.4 RQ4: General Metrics and Manual Analysis of Questions referring ChatGPT
@@
Based on the number of users active before and after the release, we perform further analysis regarding their accounts (creation date). For that, we consider active users to be the ones who were involved in posting questions, answers, and comments. Next, based on the different types of users we observe, we explore how users appeal for support on Stack Overflow, regarding questions previously asked to ChatGPT. This way, we mine all questions that refer to ChatGPT and its variated terminologies on the question’s title and body (chat-gpt, chat gpt, and gpt). [^10] We did the same for LLaMA-2, but we did not observe questions targeting it. Next, we split these questions into two groups considering their questioners’ profiles. Based on the creation date of questioners’ accounts, we have the set of questions asked by users that were reminiscent or new ones on Stack Overflow after the release of ChatGPT (reminiscent and new users). By splitting users into these categories, we expect to further explore and compare the challenges observed by them.
3.3.5 RQ5: Manual Analysis of Questions referring ChatGPT
From the questions mined in the RQ1, we filter the ones related to the domains that represent challenges for LLMs based on our findings (programming languages, and frameworks and libraries); for that, we check their associated tags, before and after ChatGPT’s release. For each domain, we check the top 10 most cited tags before ChatGPT, and then we statistically evaluate the number of questions posted addressing these different topics through a Wilcoxon test (Mann and Whitney, 1947).

### v2

@@ 3.1 Mining SO Data
To investigate the current impact of ChatGPT on Stack Overflow usage, following previous studies’ methodological steps [26], we mined SO posts and collected the data using SO’s API [^4] https://data.stackexchange.com/stackoverflow/query/new from 5 months before and after the ChatGPT release (1st July till 29th November 2022, and 30th November 2022 till 30th April 2023, respectively). Knowing that adopting the release date of ChatGPT as a cutoff may introduce bias in our results, we also mine posts dated one year after the release (30th November 2023 till 30th April 2024, respectively). This way, we can evaluate the long-term impact, considering immediate reactions and the more sustained engagement over time. This approach ensures a more comprehensive analysis, capturing shifts in discussions, challenges, or resolutions that might emerge only after extended use or broader adoption. From now on, when we say data after ChatGPT’s release, we mean the data posted just after the release. To refer to the data posted after one year of the release, we say one-year milestone. Stack Overflow’s API limits the returned data for a single request to 50 0005000050\,00050 000 rows. Hence, we performed as many requests as needed to gather all relevant data. [^5] Stack Overflow offers its archived dump dataset quarterly. By the time of our mining, we needed up-to-date information, leading us to use the provided API. For example, to get the posts posted between November 29-30th, we used the following query:
@@
One-Year Milestone
Questions 674 425674425674\,425674 425 578 115578115578\,115578 115 296 718296718296\,718296 718
Questions with Answers 424 808424808424\,808424 808 329 412329412329\,412329 412 167 303167303167\,303167 303
@@
212 775212775212\,775212 775 157 187157187157\,187157 187 78 2627826278\,26278 262
Questions without Answers 249 617249617249\,617249 617 248 703248703248\,703248 703 129 415129415129\,415129 415
Answers 543 533543533543\,533543 533 425 391425391425\,391425 391 202 326202326202\,326202 326
Comments 1 673 90616739061\,673\,9061 673 906 1 348 98213489821\,348\,9821 348 982 718 496718496718\,496718 496
Tags 36 8373683736\,83736 837 37 3283732837\,32837 328 31 8673186731\,86731 867
@@
Initially, we collected the posts and their related comments (as returned by Stack Overflow’ API). However, in Stack Overflow’s data, questions and answers are interchangeably treated as regular posts, differentiating from each other based on their type. Furthermore, while questions can receive multiple answers, only one can be targeted as the accepted one; posts can be associated with multiple comments. We collected all information available for each element under investigation, including the creation and last modification date, title, description, score, views, etc. The creation and modification dates are essential information since a question might have been added before the ChatGPT release, but its accepted answer might have been dated after ChatGPT’s release. Next, we organize the data by associating the questions with their related answers and comments. This step is required since answers and comments for questions not dated from the interval evaluated here might introduce bias in our results. For example, a question posted before the release of ChatGPT may be answered by a user that used the LLM to address the initial reported issue.
This way, we adopt one particular constraint to investigate the overall impact of posting activity on Stack Overflow. For the questions posted before ChatGPT release, we only consider valid answers and comments, the ones also dated before the release date (30th November 2022). Finally, we were able to collect 674 425674425674\,425674 425, 578 115578115578\,115578 115 and 296 718296718296\,718296 718 questions before, after, and one-year milestone ChatGPT’s release, respectively. Table 1 presents an overview of our dataset. Overall, we notice a drop of almost 100 000100000100\,000100 000 between the before and after the release of ChatGPT (for the same amount of time, i.e., five months). One year after its release, the drop becomes even more pronounced, decreasing by almost threefold (approximately 300 000300000300\,000300 000 questions).
@@
On the other hand, to investigate the impact of questions addressing specific topics (programming languages, frameworks, and libraries), we first properly filtered the questions focusing on these topics. When users ask about these topics, they commonly add the name of the target topic as one of the question tags; for example, Java, Angular, and Pandas. Since multiple tags can be used for a single question, we counted the number of times a single tag was adopted. Then, we organize the frequency of questions associated with each topic daily. Although users may consider variations or abbreviations when targeting their subjects, we observe a consistent adoption of the official topic names; for example, Pandas can be abbreviated to pd.
@@
Finally, to investigate the impact on the user’s activity patterns in Stack Overflow, we evaluate the user information associated with the final set of questions, answers, and comments collected previously. For that, we group the users associated with each content into three groups: questioners, respondents, and commentators. This way, we could compare the number of users involved in each type of activity in Stack Overflow. To answer RQ4, we collected further user information, including their account creation date, allowing us to identify reminiscent and new users after the release. We request each user’s information using their IDs on the Stack Overflow API.
@@ 3.2 Prompting LLM Models
This study focuses on two LLMs: ChatGPT 3.5 and LLaMA-2-7b. We used the GPT-3.5-turbo variant for our research, interacting with it through its API. Unlike ChatGPT, we call LLaMA by locally loading it. For our experiments, we used the 7B-sized LLaMA-2 model (Llama-2-7b-chat-hf checkpoint in its base configuration), which is the 7-billion-parameter chat-optimized variant, ensuring alignment with conversational tasks. The model was loaded via Hugging Face support [^7] https://huggingface.co/meta-llama/Llama-2-7b, on a machine equipped with 40GB RAM. When prompting the LLMs, we follow a standardized approach that simulates a prior conversation to establish context. For that, we consider the tags associated with the questions as the foundation for instructing the LLM to adopt the persona of an expert with extensive knowledge in the relevant subject matter[42]. Next, we consider the question’s title to properly ask the LLM to explain how to fix the problem. Additionally, we argue that further context will be given by informing the question description. Figure 2 presents the adopted approach with more details. Regarding the usage of these models, we consider their default configurations. For example, we prompt the models with their default temperature and context length (maximum token limit) [43]. By making this decision, we aim to avoid arbitrary bias, considering the diversity of tasks addressed in this study. Here, we focus on evaluating the models in similar ways users would, such as respecting their default configurations and interacting with them under typical usage scenarios without additional hyperparameter adjustments. Furthermore, related studies did not report specific hyperparameters, leading us to believe they used default configurations as well. Finally, for each question, we prompt each model once.
@@
Figure 2: Prompt approach adopted to prompt the LLMs. Question contents were given as input, replacing the elements informed highlighted with << and >>, respectively.
3.3 Research Questions: Evaluation Design
@@
This section presents the methodology adopted to address each RQ and the associated metrics adopted when running our study.
3.3.1 RQ1: Comparing Human and LLM-Generated Answers for Stack Overflow Questions
Textual Similarity: To assess the similarity between answers from Stack Overflow and those generated by LLMs, we employed the cosine similarity metric ($\theta$)  [44, 45]. To compute this metric, we relied on a pre-trained Sentence Transformers model (all-MiniLM-L6-v2) [^8] https://github.com/UKPLab/sentence-transformers. First, we take the accepted and LLM-generated answers for a given question and compute their embeddings. Second, we used the function cos_sim from the same model, passing the previous embeddings and computing the metric.
Semantic Similarity: Knowing that the cosine similarity is sensitive and does not compute the semantics between two given answers, we might misclassify similar answers due to low textual similarity. Motivated by previous studies [12], we decided to compute the semantic similarity relying on the judgment capability of the LLMs evaluated in this study. Our goal is not to evaluate the reliability of a particular semantic similarity metric, but rather to analyze how the LLMs themselves assess semantic similarity between answers. Specifically, we investigate the degree of consistency in their judgments and how closely they align with human evaluation. First, we simulate a prior conversation to establish the context, using the questions’ tags again to instruct the LLM to adopt the persona of an expert. Second, we ask the LLM to compare the original and LLM-generated answers, reporting an evaluation using a VERY LOW to VERY HIGH scale. By combining both answers in one single prompt, we reached the maximum number of tokens by the LLMs; ChatGPT did not classify one answer, while LLaMA did not classify 72 answers due to resource constraints, resulting in 1,464 answers that could be analyzed. Third, to evaluate the consistency and level of agreement among different runs for the same model, we prompted ChatGPT to classify the 1,535 answers three times. In the end, we randomly selected a subsample of 91 questions from each group (ensuring a 95% confidence level with a 10% margin of error). The primary author manually labeled these questions, following the same scale previously adopted by the LLMs, by comparing the contents of the human-reported and LLM-generated answers. Since the categories adopted present an ordinal structure, we must consider such aspect when assessing the level of inter-rater agreement between our manual analysis and the assessments provided by LLMs, and the agreements between the LLMs themselves. As categories have a defined order, the distance between them may not be equal; for example, disagreeing between VERY HIGH and HIGH is less severe than between VERY HIGH and VERY LOW. As a result, we compute Krippendorff’s alpha coefficient  [46] with an ordinal distance function, as this coefficient handles ordinal weighting by default.
@@ 3.3.1 RQ1: Comparing Human and LLM-Generated Answers for Stack Overflow Questions
Sentiment Analysis. In light of research suggesting that avoiding negative comments can improve answer acceptance on Stack Overflow [39], we aim to investigate whether LLMs’ responses align with this observation. To do so, we perform a sentiment analysis on answers generated by the LLMs using a pre-trained Transformers model (sentiment-analysis) [^9] https://github.com/huggingface/transformers. Subsequently, we statistically compared the sentiment outcome achieved by each LLM using the Wilcoxon test.
3.3.2 RQ2: Extracting Meta Information and Associated Domains from Questions
To answer this RQ, we perform a manual analysis focusing on the questions linked to answers reporting low and high textual similarity for each LLM evaluated. For that, we select the questions associated with the first and fourth quartile of each LLM violin plot (low and high similarity, respectively; we explain in detail when discussing our results in Section 4). Since the main focus here is to report the domains that challenge LLMs, we adopt the taxonomy provided by Barua et al. [28] as a starting point for our categorization. When analyzing the proposed categories, we observe that they were too specific, limiting their application to our sample. This way, we decided to group similar categories into single ones. For example, Barua et al. [28] report UI Development, Website Design/CSS, Web Development, and Web Service/Application as individual categories. For our study, we group all of them into Web Development, as they share a similar goal. As a result, we define nine categories: Database/SQL, Framework/Library, General Programming Concepts, Mobile, Networking, Operating Systems, Programming Languages, Tools/IDEs, and Web Development. However, when performing the analysis, we observe that some questions could not be classified using the previous categories. By analysing them, we additionally report two categories: DevOps and ML/Datascience. Finally, our analysis is based on 11 categories, nine derived from previous work and two new ones reported by us.
Aiming to get more understanding of these questions, we decided to further explore them based on the meta information we could extract. Initially, we explored some questions to identify the information shared among them. We adopted a simple process, by randomly selecting questions previously prompted to LLMs (no more than 20 questions). As our main goal was to specify the types of information shared by users when posting their questions, when we observed that no new type of information was reported (saturation), we stopped the exploration. As a result, we defined a spreadsheet, incorporating the information we could extract and their predefined values, providing a guide for the manual analysis. Table 2 provides the information we extracted. For this analysis, we focus only on the aspects associated with the questions.
Similarity Correlation. To explore possible factors contributing to the higher reliability of answers generated by LLMs, we extended our analysis to include the context in which the models were prompted. Overall, we performed a correlation analysis between (i) the textual similarity achieved by a given LLM answer, previously computed, and (ii) the number of words associated with the different types of context we have. Since we have three different types of context in our study (tags, title, and description), we adopted distinct strategies for each, as reported below. For tags, we computed the number of tags associated with each question, as this directly reflects the context provided by the tags. For both the title and description, we first computed the number of words. Then, we split them into intervals: a 2-word interval for the title and a 25-word interval for the description. The rationale behind this segmentation was to capture the varying levels of context provided by different text lengths. For titles, a 2-word interval was chosen because we did not observe any questions with only a single word. For descriptions, we opted for 25-word intervals, as there was a consistent rise in the number of questions at each interval length (e.g., 25, 50, etc.). Our goal with this analysis was to quantify the impact of context size on model performance, revealing how the amount of input text influences the quality of the LLM’s responses. Since our dataset was not drawn from a normally distributed population, as confirmed by a Shapiro-Wilk test, we used the non-parametric Spearman test. Initially, we verify whether the context used to prompt the LLMs correlates with the textual similarity level previously computed (for that, we rely on the Spearman test).
Structural Aspects Definition Associated Values Target RQs
Question Context Adequacy
The level of information
provided in the question.
Shallow, Medium, Strong RQ2, RQ4
External Reference
External content linked
with the question
Another SO question,
documentation, link
No Textual Content
Adoption of content
different than text
Image
Illustrative Context
Supportive information
provided to illustrate the
target problem
Code, Error Message,
Outcome, Expected Outcome
Goal
The approach adopted
to ask for assistance
Question, Explanation,
Open
Domain
The domain associated
with the question
Categories adapted from [28]
Tag Concordance
The level of concordance
among the informed tags
Shallow, Medium, Strong
Answer No Textual Content
Adoption of content
different than text
Image RQ4
External Reference
External content linked
with the answer
Another SO question,
documentation, link
Illustrative Content
Supportive information
provided to illustrate
the solution
Code, Outcome,
Expected Outcome
Table 2: Structural aspects extracted from Stack Overflow questions and associated accepted answers. For a given aspect, we present its definition and possible associated values.
3.3.3 RQ3: Analyzing General and Domain-Specific Impact of Posting Activity on Stack Overflow
To assess the overall impact of ChatGPT’s release on question, answer, and comment frequency, we grouped the frequency of each element by the day they were posted and then performed our analysis. To evaluate the impact on domains, we adopt a similar approach by grouping the questions based on the challenging domains reported in RQ2 (programming languages, and frameworks and libraries). For each domain, we manually searched for the top 10 most cited tags, which were the same before and after ChatGPT’s release. Then, for each tag, we compute the frequency at which they were used daily. Although Frameworks and Libraries are classified as one single domain, we explore them individually, selecting the ten most cited frameworks and libraries.
Knowing that our data did not follow a normal distribution, as verified by the Shapiro-Wilk test [47],101010p-value < 0.01 we employed the non-parametric Wilcoxon test to compare the distributions of each evaluated group, using a significance level of $\alpha=0.05$ [48].
3.3.4 RQ4: Checking Impact on User’s Activity Patterns in Stack Overflow
Based on the number of users active before and after the release of ChatGPT, we first investigate the impact on the different types of users (questioners, respondents, and commentators). Different from the previous RQ, in this one, we focused only on the data before and after release. Knowing that our primary goal is to evaluate the impact of ChatGPT’s release on user activity, to ensure a focused comparison, we analyzed data from five months before and up to one year after its release (after and one-year milestone). While we acknowledge that external factors could influence user activity both before and after the release, restricting the pre-ChatGPT period helps minimize historical biases unrelated to ChatGPT, such as long-term platform trends, policy changes, technologies released, or shifts in users’ activity patterns. While post-release activity may also be affected by other factors, by adopting a one-year window, we believe we can ensure that we analyze both immediate and sustained changes, making it more likely that observed trends are linked to ChatGPT’s introduction rather than unrelated fluctuations.
@@ 3.3.4 RQ4: Checking Impact on User’s Activity Patterns in Stack Overflow
Moving on, we first compute the IDs of the users associated with the questions and then compare which users were active or inactive after ChatGPT’s release. Second, we further compare the active users after release based on the creation date of their accounts. Third, based on the different types of users we observe (reminiscent and new users), we explore how users appeal for support on Stack Overflow, regarding questions previously asked to ChatGPT. This way, we mine all questions that refer to ChatGPT and its varying terminologies on the question’s title and body (chat-gpt, chat gpt, and gpt). [^11] We did the same for LLaMA-2, but we did not observe questions targeting it.
Initially, by analyzing all the 578 115578115578\,115578 115 questions after ChatGPT release (see Table 1), which also include questions without answers, we observe that 2193 questions mention ChatGPT in their contents (title or body). Next, checking the creation date of the questioners’ accounts, we observe that $1439$ questions were asked by reminiscent users, while 731 questions were asked by new ones. The remaining questions were asked anonymously or by users no longer with active accounts in Stack Overflow, so we could not check whether the questions were associated with reminiscent or new users. For each group, we randomly select a subsample of questions to be manually analyzed (91 and 86 questions, respectively), aiming for a 95% confidence level and a 10% margin of error. Then, we manually analyze each question, extracting the information reported in Table 2. For the current analysis, we focus on the question and the associated accepted answer, as we are also interested in how the respondents would behave when facing a problem that LLMs were previously asked.